OpenStudy (anonymous):

what is the simplified form of (1/x)-2/x^2+x

jimthompson5910 (jim_thompson5910):

is that \[\Large \frac{1}{x} - \frac{2}{x^2+x}\] ??

OpenStudy (anonymous):

you know how to do it

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok in order to subtract, the denominators must be the same

jimthompson5910 (jim_thompson5910):

so you need to get both denominators equal to the LCD

jimthompson5910 (jim_thompson5910):

in this case, the LCD is x(x+1) so how do you get the first denominator equal to the LCD?

jimthompson5910 (jim_thompson5910):

you have x as the denominator of the first fraction you want x(x+1) so what's missing?

OpenStudy (anonymous):

idunno

jimthompson5910 (jim_thompson5910):

You're missing an (x+1), so you multiply top and bottom of the first fraction by this to get \[\Large \frac{1}{x} - \frac{2}{x^2+x}\] \[\Large \frac{1(x+1)}{x(x+1)} - \frac{2}{x^2+x}\] \[\Large \frac{x+1}{x(x+1)} - \frac{2}{x^2+x}\]

jimthompson5910 (jim_thompson5910):

notice how x(x+1) distributes to x^2 + x, so we can further say \[\Large \frac{x+1}{x(x+1)} - \frac{2}{x^2+x}\] turns into \[\Large \frac{x+1}{x^2+x} - \frac{2}{x^2+x}\]

OpenStudy (anonymous):

so what am i missing

jimthompson5910 (jim_thompson5910):

now you just need to subtract the numerators

jimthompson5910 (jim_thompson5910):

and place that difference over the common denominator

OpenStudy (anonymous):

ok so what is that

jimthompson5910 (jim_thompson5910):

x+1 - 2 = ???

OpenStudy (anonymous):

=0

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

ok i dunnno get it

jimthompson5910 (jim_thompson5910):

1 - 2 = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so x + 2 - 1 turns into x - 1

jimthompson5910 (jim_thompson5910):

and \[\Large \frac{x+1}{x^2+x} - \frac{2}{x^2+x}\] turns into \[\Large \frac{x-1}{x^2+x}\]

jimthompson5910 (jim_thompson5910):

which is your final answer

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{1}{x} - \frac{2}{x^2+x}\] fully simplifies to \[\Large \frac{x-1}{x^2+x}\]

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):